OpenStudy (bunnyzrockgirl):

if i get 5 questions out of 8 questions right, do i have an 80% or above?

OpenStudy (bunnyzrockgirl):

@boricua5 @Abhisar @campbell_st anyone?

OpenStudy (campbell_st):

no... 5/8 is between 4/8 = 50% and 6/8 = 75% so the key is 1/8 = 12.5% multiply by 5 to get 5/8 as a percentage

OpenStudy (bunnyzrockgirl):

so in order for me to get at least an 80 on my worksheet, i have to get 7 out of the 8 right?

OpenStudy (campbell_st):

well that's correct... thay would be more than 80%

OpenStudy (bunnyzrockgirl):

thx, i cant seem to get the last 3 answers on my paper... no one can answer my problems... :(